During the past five hours, the Port of Long Beach in California unloaded large quantities of grain from ships and relevant data are summarized in the table below:

Hour Quantity unloaded (tons)

--------------------------------------------

1 3,406

2 3,440

3 3,298

4 3,372

5 3,436

--------------------------------------------

1) Use the naïve approach to develop a forecast of the tonnage of grain unloaded in each hour up to the 5th hour.

(2) Repeat (1) above by using the 2-hour simple moving average method.

(3) Which of the two forecasting techniques is more accurate based on MAD? Be sure to carry four decimal places for irrational numbers.
Answer #1

1) Naive Approach

F(2) = 3406

F(3) = 3440

F(4) = 3298

F(5) = 3372

2. 2-hour moving average

F(3) = (3406+3440)/2

= 3423

F(4) = (3440+3298)/2

= 3369

F(5) = (3298+3372)/2

= 3335

3. MAD for Naive

MAD = (34 + 142 + 74 + 64)/4

= 78.5

MAD for Moving Average

MAD = (125 + 3 + 101)/3

= 76.3333

Hence the 2 - hour simple.moving average is more accurate